OpenStudy (sarahc):

Eliminate the parameter. x = t2 + 1, y = t2 - 1 y = x + 2, x ≥ 1 y = x - 2, x ≥ 1 y = x2 - 2, x ≥ 1 y = x2 + 2, x ≥ 1

OpenStudy (jdoe0001):

\(\bf x=t^2+1\qquad y=t^2-1 \\ \quad \\ x=t^2+1\impliedby \textit{solve for "t", what do you get?}\)